OpenStudy (anonymous):

how do you find the line of symmetry for the equation: 24x-16x^2?

OpenStudy (shiraz14):

This is a parabola with a max turning point. Do you think you can find the coordinates of the vertex?

OpenStudy (anonymous):

yeah

OpenStudy (tkhunny):

\(24x - 16x^{2} = 8(3x-2x^{2}) = 8x(3-2x)\) - Rather staring you in the face from there.

OpenStudy (anonymous):

0,1.5?

OpenStudy (shiraz14):

@Erob : Those are the roots of the equation - the parabola is symmetric about its vertex (& the x-coordinate of the vertex is at the mid-point of the x-coordinate of its roots) - so what do you think the equation for the line of symmetry is?

OpenStudy (anonymous):

the equation is \[x=-\frac{ b }{ 2a }\] but I'm having trouble filling in the blanks

OpenStudy (anonymous):

does b=24, a=-16?

OpenStudy (shiraz14):

Yes, you can do that, or you can find the mid-point of the x-coordinates of the roots. In either case, the answer is the same, i.e. x=0.75.

OpenStudy (anonymous):

ok I just wanted to check because my teacher doesn't usually give question where the answer is a decimal. thanks for the help!

OpenStudy (shiraz14):

sure, you're welcome